1. Write an equation of the line containing the given point and parallel to the given line.
(5,7); x+7y=8
The equation of the line is y=__________
Express your answer in the form y=mx+b.
2. Write an equation of the line containing the given point and parallel to the given line.
(-5,2); 4x=5y+7
The equation of the line is y=__________
Express your answer in the form y=mx+b.
3.Write an equation of the line containing the given point and perpendicular to the given line.
(4,8); 9x+y=8
The equation of the line is y=__________
Express your answer in the form y=mx+b.
I really tried to figure this out but I am running out of time and I cannot figure them out. Please help me, thank you.

Answer by Alan3354(30955)

1. Write an equation of the line containing the given point and parallel to the given line.
(5,7); x+7y=8
Find the slope of the line by solving for y
x+7y = 8
y = (-1/7)x + 8/7
Slope m = -1/7
-----------------
Use y = mx + b and the point to find b
7 = (-1/7)*5 + b
b = 54/7
y = (-1/7)x + 54/7
--------------------
2. Write an equation of the line containing the given point and parallel to the given line.
(-5,2); 4x=5y+7
Do it like #1
-------------------
3.Write an equation of the line containing the given point and perpendicular to the given line.
(4,8); 9x+y=8
Find the slope
y = -9x + 8
m = -9
The slope of lines perpendicular is the negative inverse, +1/9
Use y = mx + b and the point as in #1
-----------------------------